In the month of May, local highschool students sent soldiers 289 letters. They sent about the same number of letters every day of the month. What os the best estimate of the number of letters that were sent each day?
1
Expert's answer
2012-12-06T09:10:42-0500
The number of letters that were sent each day is n = T / D,where T is the total number of letters sent, and D is the number of days. There
are 31 days in May.
n = 289 / 31 = 9.323
Answer: approx. 9 letters per day.